OpenStudy (anonymous):

completely factor w^2-17w+42

OpenStudy (anonymous):

find two numbers that when multiply gives you 42, and when added gives you -17w

OpenStudy (anonymous):

or use quadratic formula

OpenStudy (anonymous):

(w-14) (w-3) -14 * -3 = 42 and -14 + -3 = -17

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\] \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (unklerhaukus):

\[(x-x_1)(x-x_2)=0\]